Entries for Stock Dividends

Healthy Life Co. is an HMO for businesses in the Fresno area. The following account balances appear on Healthy Life’s balance sheet: Common stock (400,000 shares authorized ; 5,000 shares issued), $100 par, $500,000; Paid-In Capital in excess of par—common stock, $100,000; and Retained earnings, $4,000,000. The board of directors declared a 1% stock dividend when the market price of the stock was $121 a share. Healthy Life reported no income or loss for the current year.

If no entry is required, select "No entry required" and leave the amount boxes blank. If an amount box does not require an entry, leave it blank.

a1. Journalize the entry to record the declaration of the dividend, capitalizing an amount equal to market value.

Stock Dividends
Stock Dividends Distributable
Paid-In Capital in Excess of Par-Common Stock

Feedback

a1. Recall that a stock dividend affects only stockholders' equity.

a2. Journalize the entry to record the issuance of the stock certificates.

Stock Dividends Distributable
Common Stock

Feedback

b. Determine the following amounts before the stock dividend was declared: (1) total paid-in capital, (2) total retained earnings, and (3) total stockholders' equity.

Total paid-in capital $
Total retained earnings $
Total stockholders' equity $

c. Determine the following amounts after the stock dividend was declared and closing entries were recorded at the end of the year: (1) total paid-in capital, (2) total retained earnings, and (3) total stockholders' equity.

Total paid-in capital $
Total retained earnings $
Total stockholders' equity $
Answer #1

a1. Journalize the entry to record the declaration of the dividend, capitalizing an amount equal to market value.

Stock Dividends (5000*1%*121) 6050
Stock Dividends Distributable 5000
Paid-In Capital in Excess of Par-Common Stock 1050

Feedback

a1. Recall that a stock dividend affects only stockholders' equity.

a2. Journalize the entry to record the issuance of the stock certificates.

Stock Dividends Distributable 5000
Common Stock 5000

Feedback

b. Determine the following amounts before the stock dividend was declared: (1) total paid-in capital, (2) total retained earnings, and (3) total stockholders' equity.

Total paid-in capital $500000+100000 = 600000
Total retained earnings $4000000
Total stockholders' equity $4600000

c. Determine the following amounts after the stock dividend was declared and closing entries were recorded at the end of the year: (1) total paid-in capital, (2) total retained earnings, and (3) total stockholders' equity.

Total paid-in capital $600000+6050 = 606050
Total retained earnings $4000000-6050 = 3993950
Total stockholders' equity $4600000